A survey of 1000 students at a large university asked students about their view on capital punishment. Half of the students surveyed were men and half were women. Of the men surveyed, 280 were opposed to capital punishment. Of the women surveyed, 320 were opposed.
How many students in the sample were in favor of capital punishment?
How many of those in favor of capital punishment were women?

Found 2 solutions by greenestamps, math_tutor2020:
Answer by greenestamps(13200)



You only need to be able to read (and understand) and do very simple addition and subtraction to answer the questions....

1000 students, half men and half women, so 500 men and 500 women.

280 men opposed to capital punishment, so 500-280 in favor.

320 women opposed to capital punishment, so 500-320 in favor.

Do those calculations to find the answers. You shouldn't need our help to do that....


Answer by math_tutor2020(3817)


We can organize the data into a two-way table.
MenWomenTotal
In favorABC
Opposed280320D
Total5005001000

I'll let the student fill in the gaps.

Hopefully it is fairly obvious that we have the following equations when looking at each column separately
A+280 = 500
B+320 = 500
C+D = 1000

We also have these three equations when focusing on the rows separately
A+B = C
280+320 = D
500+500 = 1000

Another equation that can be formed is this
A+B+280+320 = 1000
which represents adding up the inner 2x2 block of values to get the grand total